OpenStudy (anonymous):

How would I find the common multiples of 20, 44 and 88

OpenStudy (mathhelp346):

2

OpenStudy (anonymous):

actually 2 is a common FACTOR multiples are multiples, as in 20,40,60,80,...

OpenStudy (anonymous):

\[20-2^2\times 5\] \[44=2^2\times 11\] \[88=2^3\times 11\] so one multiple will be \[2^3\times 5\times 11=440\]